a cyclist of mass 82 kg is riding his bicycle of mass 13 kg up a 1 in 10 incline. What power does he need to develop in other to maintain aconstant speed of 24kph

1
Expert's answer
2022-02-17T10:05:39-0500

Given:

"m=\rm 82\: kg+13\: kg=95\: kg"

"\sin\theta=1/10"

"v=24\:\rm km/hr=6.67\: m/s"


The net force



"F=mg\sin\theta=95*9.8*1/10=93.1\:\rm N"

The total work done



"W=Fd=Fvt"


The required power


"P=\frac{W}{t}=Fv=93.1*6.67=621\:\rm W"